Care and Feeding May 27, 2007

Physics 214 General Physics II

The Care and Feeding of a Physics Problem Set

A Checklist for Students

1. Start with a basic equation. In each chapter, I will indicate which are the basic equations. These equations represent fundamental physics principles. You are expected to start with these equations; simply pulling an equation from an example will not receive full credit.

2. Explain your work clearly, in English. Do not simply repeat in words what your equations say. Instead, explain why the equation you chose is relevant to the problem at hand. If you need to combine several equations to solve a problem, explain the relationships between them or how you are combining them.

3. Have a high E/M (English/Math) ratio. This is the same as rule number 2, but is so important it’s worth repeating.

4. Indicate your assumptions explicitly. What physical approximations are you using? Is the problem open to interpretation? If so, explain your interpretation.

5. Define your variables. Don’t assume we know what you mean with each variable. When in doubt, define.

6. Include proper diagrams whenever possible, especially free-body diagrams. Don’t underestimate how much drawing a diagram can clarify the problem. For any problem where you apply Newton’s laws, you must draw a free-body diagram, or you will receive half credit.

7. Solve the problem symbolically. You must arrive at an equation of the form “desired quantity = expression in terms of other variables” before plugging in numbers. If you do not, you will receive half-credit.

8. Use a reasonable number of digits. No matter how many digits your calculator gives you, your answer must have significant digits consistent with the precision of the numbers in the problem.

9. Apply the “Gravity Does Not Point Up” test. Think about your final answer. If it is completely unreasonable, rework the problem or at least explain why it cannot be right (you may receive additional partial credit for doing so).

10. Get your units right. Units are easy points and serve as a check that you answered the correct question.

Care and Feeding May 27, 2007

Physics 214 General Physics II

Copyright Notice

This document is copyrighted \begin{matrix}c \\ ◯ \end{matrix} 2004-7 by Itai Seggev. It may, however, be freely redistributed or used in derivative works for noncommercial purposes and with proper attribution under the terms of the Creative Commons Attribution-NonCommercial-ShareAlike License, Version 3.0. This license should be held to apply equally to all forms of this document, whether print or electronic, including but not limited to any current or future file format. However, the requirements (i) that the full copyright notice be included; (ii) that the disclaimer of warranty be included; and (iii) that the license text or its URI be included; as detailed in sections 4(a) and 4(d) of the license, are hereby waived by the author provided all of the following conditions are met:

1. the copies are distributed exclusively to students enrolled in a course—which is taught by the licensee—at a secondary or post-secondary school;

2. the copies are distributed in hard copy or binary-only form (e.g., PDF); if distributed in binary form via a computer network, the URI of the copy must be restricted to students enrolled in the course;

3. the author is credited in the following or some substantially similar form: “Written by Itai Seggev” for copies of the Work and “Based on a Work by Itai Seggev” for derivatives of the Work.

For details of the license see http://www.creativecommons.org/licenses/by-nc-sa/3.0/.